# Optimal Fairness under Local Differential Privacy

### v1

@@ V Experiments
The optimal mechanisms introduced in Section IV are designed to minimize data unfairness while satisfying utility and LDP constraints. We refer to both the binary and non-binary mechanisms as OPT. In this section, we empirically evaluate the effectiveness of OPT across different classification settings. Section V-A compares OPT with existing LDP mechanisms for both binary and multi-valued sensitive attributes. Section V-B evaluates OPT against RR under the fairness framework of [mozannar2020fair]. Finally, Section V-C compares OPT with Fair Projection [alghamdi2022beyond] and FairBalance [yu2024fairbalance], two state-of-the-art post- and pre-processing fairness methods [wang2024aleatoric, soremekun2022software]. [^1] Experimental code is publicly available at https://github.com/hradghoukasian/ldp_fairness.
@@ V-A Comparison with Existing LDP Mechanisms
Following [arcolezi2025group], we evaluate the proposed mechanism on the Adult [UCI] and Law School Admissions Council (LSAC) [wightman1998lsac] datasets, where the label $Y$ indicates income above \$26k and bar-exam pass, respectively. Depending on the setting (binary or multi-valued), the sensitive attribute is chosen among race, gender, race--gender, or family income.
@@
We next evaluate non-binary attributes: race ($k=5$) and race--gender ($k=10$) for Adult, and family income ($k=5$) for LSAC. The optimal mechanism is computed by solving the min–max linear fractional program in (9) following [jiao2022solving], using the largest feasible $\zeta$ for each $\varepsilon$. Results are compared against GRR, SS, and the non-private baseline.
@@
V-B Comparison under the [mozannar2020fair] Framework
We integrate OPT into the two-step post-processing framework of [mozannar2020fair], which is designed to train fair classifiers when only privatized sensitive attributes are available. In their original setup, RR is used to ensure LDP; we replace RR with OPT. The procedure first trains a base classifier using the privatized attributes, then post-processes its predictions to enforce equalized odds. To ensure a fair comparison and demonstrate the adaptability of OPT, we apply OPT in both the pre-processing of sensitive attributes and the post-processing step.
@@ V-B Comparison under the [mozannar2020fair] Framework
Figure 6: Comparison of the two-step framework of [mozannar2020fair] using RR and OPT.
Figure 6 shows that on LSAC, OPT and RR achieve similar accuracy, while OPT yields smaller $\Delta_{\mathrm{EOd}}$ across most $\varepsilon$ values. On Adult, OPT attains lower $\Delta_{\mathrm{EOd}}$ for larger $\varepsilon$ with comparable accuracy, and higher accuracy for smaller $\varepsilon$ at the cost of increased unfairness. These results indicate that OPT can effectively replace RR in the [mozannar2020fair] framework.
@@ V-C Comparison with Fairness Pre/Post-Processing Methods
FairProjection. FairProjection is a post-processing fairness method that maps a pre-trained, potentially unfair classifier onto the space of models satisfying a chosen group-fairness criterion. The projected model is constructed by post-processing the predictions of the original classifier to enforce the target fairness metric. Following [alghamdi2022beyond], we compare OPT and FairProjection on the Adult and COMPAS [angwin2022machine] datasets. In Adult, the label $Y$ and the sensitive attribute are defined as before; in COMPAS, $Y$ denotes recidivism with sensitive attribute race. Both methods are evaluated under mean equalized odds and statistical parity constraints using gradient boosting. Models use a 70/30 train–test split, averaged over 30 trials with 95% confidence intervals. FairProjection provides a controllable fairness–accuracy trade-off by allowing one to specify a target unfairness gap to be satisfied. To enable a fair comparison with OPT, we evaluate OPT at $\varepsilon=1$ and compare it to FairProjection at matched accuracy, reporting the corresponding $\Delta_{\mathrm{SP}}$ and $\Delta_{\mathrm{MEO}}$ values (Table I).
FairBalance. FairBalance is a pre-processing bias mitigation algorithm that balances the class distribution within each demographic group by assigning appropriate weights to training samples. Its goal is to identify the underlying causes of fairness violations and mitigate them without altering the standard training procedure. Following [yu2024fairbalance], we evaluate OPT and FairBalance on the Bank Marketing, Student Performance (Math), and German Credit datasets [UCI]. In Bank, $Y$ indicates term-deposit subscription and the sensitive attribute is age; in Math, $Y$ denotes final grade above a threshold and the sensitive attribute is gender; in German, $Y$ indicates credit risk and the sensitive attribute is gender. All models use logistic regression with 70/30 train–test splits, averaged over 30 trials and 95% confidence intervals. We fix $\varepsilon=1$ and compare OPT to FairBalance, reporting $\Delta_{\mathrm{EO}}$ and $\Delta_{\mathrm{MEO}}$ (Table II).
@@ VI Conclusion
Motivated by recent empirical findings on the impact of local differential privacy LDP on fairness in classification, this work introduces an optimal LDP mechanism applied to sensitive attributes to minimize data unfairness. For binary $A$, we derive a closed-form expression for the optimal mechanism; for multi-valued $A$, we provide an equivalent formulation solvable numerically. We further show that when DA–optimal classifiers in $\mathcal{H}^{*}$ are trained on data with reduced unfairness, the resulting classifiers also exhibit improved fairness (measure by statistical parity). These theoretical results are supported by experiments on multiple real-world datasets. On the Adult and LSAC datasets, the proposed mechanism effectively reduces classification unfairness while maintaining utility comparable to state-of-the-art LDP mechanisms at a fixed $\varepsilon$-LDP level. Moreover, OPT can directly replace RR mechanism in the fairness framework of [mozannar2020fair], enabling fairness-aware learning when only the privatized sensitive attribute is available. Finally, our approach achieves a better accuracy–fairness trade-off than both pre- and post-processing fairness interventions, while preserving the privacy of sensitive attributes across the Adult, COMPAS, Bank, Math, and German datasets.

### v2

@@ V Experiments
The optimal mechanisms introduced in Section IV are designed to minimize data unfairness while satisfying utility and LDP constraints. We refer to both the binary and non-binary mechanisms as OPT. In this section, we empirically evaluate the effectiveness of OPT across different classification settings. Section V-A compares OPT with existing LDP mechanisms for both binary and multi-valued sensitive attributes. Section V-B evaluates OPT against RR under the fairness framework of [54]. Finally, Section V-C compares OPT with Fair Projection [4] and FairBalance [72], two state-of-the-art post- and pre-processing fairness methods [62, 61]. [^1] Experimental code is publicly available at https://github.com/hradghoukasian/ldp_fairness.
@@ V-A Comparison with Existing LDP Mechanisms
Following [7], we evaluate the proposed mechanism on the Adult [25] and Law School Admissions Council (LSAC) [68] datasets, where the label $Y$ indicates income above \$26k and bar-exam pass, respectively. Depending on the setting (binary or multi-valued), the sensitive attribute is chosen among race, gender, race--gender, or family income.
@@
We next evaluate non-binary attributes: race ($k=5$) and race--gender ($k=10$) for Adult, and family income ($k=5$) for LSAC. The optimal mechanism is computed by solving the min–max linear fractional program in (9) following [40], using the smallest feasible $\zeta$ for each $\varepsilon$. Results are compared against GRR, SS, and the non-private baseline.
s
@@
V-B Comparison under the [54] Framework
We integrate OPT into the two-step post-processing framework of [54], which is designed to train fair classifiers when only privatized sensitive attributes are available. In their original setup, RR is used to ensure LDP; we replace RR with OPT. The procedure first trains a base classifier using the privatized attributes, then post-processes its predictions to enforce equalized odds. To ensure a fair comparison and demonstrate the adaptability of OPT, we apply OPT in both the pre-processing of sensitive attributes and the post-processing step.
@@ V-B Comparison under the [54] Framework
Figure 6: Comparison of the two-step framework of [54] using RR and OPT.
Figure 6 shows that on LSAC, OPT and RR achieve similar accuracy, while OPT yields smaller $\Delta_{\mathrm{EOd}}$ across most $\varepsilon$ values. On Adult, OPT attains lower $\Delta_{\mathrm{EOd}}$ for larger $\varepsilon$ with comparable accuracy, and higher accuracy for smaller $\varepsilon$ at the cost of increased unfairness. These results indicate that OPT can effectively replace RR in the [54] framework.
@@ V-C Comparison with Fairness Pre/Post-Processing Methods
FairProjection. FairProjection is a post-processing fairness method that maps a pre-trained, potentially unfair classifier onto the space of models satisfying a chosen group-fairness criterion. The projected model is constructed by post-processing the predictions of the original classifier to enforce the target fairness metric. Following [4], we compare OPT and FairProjection on the Adult and COMPAS [5] datasets. In Adult, the label $Y$ and the sensitive attribute are defined as before; in COMPAS, $Y$ denotes recidivism with sensitive attribute race. Both methods are evaluated under mean equalized odds and statistical parity constraints using gradient boosting. Models use a 70/30 train–test split, averaged over 30 trials with 95% confidence intervals. FairProjection provides a controllable fairness–accuracy trade-off by allowing one to specify a target unfairness gap to be satisfied. To enable a fair comparison with OPT, we evaluate OPT at $\varepsilon=1$ and compare it to FairProjection at matched accuracy, reporting the corresponding $\Delta_{\mathrm{SP}}$ and $\Delta_{\mathrm{MEO}}$ values (Table I).
FairBalance. FairBalance is a pre-processing bias mitigation algorithm that balances the class distribution within each demographic group by assigning appropriate weights to training samples. Its goal is to identify the underlying causes of fairness violations and mitigate them without altering the standard training procedure. Following [72], we evaluate OPT and FairBalance on the Bank Marketing, Student Performance (Math), and German Credit datasets [25]. In Bank, $Y$ indicates term-deposit subscription and the sensitive attribute is age; in Math, $Y$ denotes final grade above a threshold and the sensitive attribute is gender; in German, $Y$ indicates credit risk and the sensitive attribute is gender. All models use logistic regression with 70/30 train–test splits, averaged over 30 trials and 95% confidence intervals. We fix $\varepsilon=1$ and compare OPT to FairBalance, reporting $\Delta_{\mathrm{EO}}$ and $\Delta_{\mathrm{MEO}}$ (Table II).
@@ VI Conclusion
Motivated by recent empirical findings on the impact of local differential privacy LDP on fairness in classification, this work introduces an optimal LDP mechanism applied to sensitive attributes to minimize data unfairness. For binary $A$, we derive a closed-form expression for the optimal mechanism; for multi-valued $A$, we provide an equivalent formulation solvable numerically. We further show that when DA–optimal classifiers in $\mathcal{H}^{*}$ are trained on data with reduced unfairness, the resulting classifiers also exhibit improved fairness (measure by statistical parity). These theoretical results are supported by experiments on multiple real-world datasets. On the Adult and LSAC datasets, the proposed mechanism effectively reduces classification unfairness while maintaining utility comparable to state-of-the-art LDP mechanisms at a fixed $\varepsilon$-LDP level. Moreover, OPT can directly replace RR mechanism in the fairness framework of [54], enabling fairness-aware learning when only the privatized sensitive attribute is available. Finally, our approach achieves a better accuracy–fairness trade-off than both pre- and post-processing fairness interventions, while preserving the privacy of sensitive attributes across the Adult, COMPAS, Bank, Math, and German datasets.